# Live Fitness Coaching as a Testbed for Situated Interaction

### v1

@@ Abstract
Tasks at the intersection of vision and language have had a profound impact in advancing the capabilities of vision-language models such as dialog-based assistants. However, models trained on existing tasks are largely limited to turn-based interactions, where each turn must be stepped (i.e., prompted) by the user. Open-ended, asynchronous interactions where an AI model may proactively deliver timely responses or feedback based on the unfolding situation in real-time are an open challenge. In this work, we present the QEVD benchmark and dataset which explores human-AI interaction in the challenging, yet controlled, real-world domain of fitness coaching – a task which intrinsically requires monitoring live user activity and providing timely feedback. It is the first benchmark that requires assistive vision-language models to recognize complex human actions, identify mistakes grounded in those actions, and provide appropriate feedback. Our experiments reveal the limitations of existing state of the art vision-language models for such asynchronous situated interactions. Motivated by this, we propose a simple end-to-end streaming baseline that can respond asynchronously to human actions with appropriate feedbacks at the appropriate time.
@@ 1 Introduction
Datasets that combine visual information and language have greatly contributed to advancing the abilities of AI models over the past years, ranging from captioning [17] to visual questions answering [4], to visual dialogue [15] and beyond. Particularly impressive recent showcases of this progress are recent models, such as GPT-4o and Gemini [52], which interact with a user in real-time.
@@
Despite the impressive recent progress, existing vision-language models still lag far behind human capabilities. State of the art models may be queried e.g., through prompting, in real-time for events taking place in the real-world, but they lack the ability to asynchronously interact with the user as demanded by the situation rather than when prompted. Such interactive scenarios that are grounded in the spatial and temporal context of an unfolding situation are commonly referred to as “situated” [3, 10, 11]. Addressing such situated interactive scenarios are key to the development of real-world assistive vision-language models.
@@
A notable type of situated interaction is the instructional or coaching scenario, where a coach or an instructor guides a user through a complex activity, e.g., live fitness coaching. The real-world domain of live fitness coaching has several benefits that make it an ideal test-bed for studying situated assistive vision-language models: First, fitness routines has a controlled structure, in that users are expected to (but may or may not) follow a specific series of actions. Second, despite its controlled structure, coaching is a highly challenging unsolved problem for current assistive vision-language systems. Third, it is a real-world application, whose solution can provide real and tangible benefits.
@@
Currently available large-scale video datasets [21, 24, 41, 49, 65] provide a rich set of annotations with expert demonstrations in domains such as cooking or house-hold activities. Expert demonstrations alone are not sufficient for real-world instructional scenarios such as live fitness coaching where the user might make mistakes. This has been addressed recently for some ego-centric instruction tasks by [8] and [56]. Successfully guiding a user through a fitness routine additionally necessitates the capability of understanding fine-grained human actions, providing appropriate instructions to complete the target fitness routine and providing corrective feedback – grounded in those fine-grained human actions – to correct mistakes made by the user.
@@
Overall, our main contributions are: (1) We propose the first large scale benchmark and dataset, Qualcomm Exercise Videos Dataset (QEVD)***https://developer.qualcomm.com/software/ai-datasets/qevd, aimed at the development of video-language models for live coaching. QEVD contains over 474+ hours of videos for fitness activity recognition and coaching. It includes short-clip videos (QEVD-FIT-300K) ($\sim$5 seconds in length) annotated with $1$M+ question-answer pairs, and long-range videos (>3 minutes in length) annotated with live feedbacks (QEVD-FIT-COACH, cf., Fig. 1); (2) We perform an extensive evaluation of state-of-the-art VLMs on this task, which shows that the task is a largely unsolved with significant room for improvement; (3) As a step towards closing the gap towards situated interaction, we propose a novel video LM, Stream-VLM which, instead of being constrained to turn-based interactions, can decide on-the-fly when and what to say to the user, including corrective feedbacks at the appropriate time. It constitutes an end-to-end trained baseline for real-time visual interaction with a user based on live camera input.
@@ 2 Related Work
Datasets for Activity Recognition. There is a large body of work on visual activity recognition. This includes NTU RGB+D 120 [45], FineGym [46], UCF101 [50], Kinetics [12], Moments in Time [42], ActivityNet [24], AVA-Kinetics [29], Charades [49], Something-Something [19], Something-Else [39] and others. While these datasets focus on a wide range of human activities, they contain only a few fitness activity related classes, if any (Tab. 1, col 2). Unlike our QEVD benchmark and dataset, these datasets do not have any detailed multi-modal annotations, such as questions and feedbacks, and they are do not include long videos with multiple actions or events. This restricts their utility in the development of interactive video-language models.
@@
Datasets for Procedural Activities. The Epic-Kitchens [14] dataset provides ego-centric videos of non-scripted daily kitchen activities, with post-hoc recorded narrations. YouCook2 [65] provides cooking videos annotated with instructions on how to prepare specific meals, largely featuring expert chefs. HowTo100M [41] provides narrated instructional videos, featuring a variety of activities. These datasets are not interactive (Tab. 1, col 4) as they feature first-person instructions or narrations. Furthermore, they largely feature experts and does not include mistakes likely to be made by novices (Tab. 1, col 5). In contrast, QEVD is interactive, features participants with diverse skill levels, thus includes mistakes likely to be made by novices.
@@
Ego-Exo4D [21] includes a highly diverse set of activities, performed by participants with a variety of skill levels. Assembly-101 [44] features videos of people with diverse skill levels assembling and disassembling 101 “take-apart” toy vehicles. However, these datasets are not interactive and they do not include corrective feedbacks in case of mistakes (Tab. 1, col 6). QEVD is interactive and includes corrective feedbacks from the perceptive of the fitness coach.
@@
On the other hand, WTAG [8] and HoloAssist [56] include corrective feedbacks and are focused at the development of interactive AI assistants. However, they focus on domains such as cooking or object manipulation. As they are recorded from an ego-centric perspective, they do not contain complex human actions (Tab. 1, col 3). Furthermore, while they include mistakes and associated corrective feedbacks, they do not include diverse examples of possible mistakes per target task (Tab. 1, col 7). QEVD is recorded from the perspective of a virtual fitness coach and includes fine-grained human actions – diverse exercises and their variations including a wide diversity of mistakes per exercise.
@@
Datasets for VQA and Reasoning. ActivityNet Captions [24], VATEX [55], TRECVID [5], HD-VILA [58], TGIF [31], WebVid [6], Charades [49], STAR [57] and AGQA [22] among others, focus largely on video captioning and question answering tasks. Finally, there exists a wide range of datasets on visual reasoning [4, 27, 43], including visual dialogue (for example, [15]), all of which, in contrast to our work, are based on still images not video.
@@
Models for Situated Interactions. There is also a growing body of work on extending LMs to generally reason over visual input [2, 23, 37, 48, 51, 54, 59, 61, 64]. However, these models can answer only high-level questions over images or videos about depicted scenes and objects. Models for exercise feedback are discussed in [18, 47] among others. However, such models are based on the recognition of whether or not an exercise was performed, or on counting repetitions, rather than providing interactive guidance and reasoning about the user’s movements from a third-person viewpoint, which is the focus of this work.
@@
Total High-level Questions 543,117 31,034 - -
Total Fine-grained Questions 402,973 27,047 - -
@@
Total Feedbacks 610,776 39,327 5,403 2,484
Average Feedbacks per Video†† 2.0 $\pm$ 10.1 2.4 $\pm$ 6.9 5.0 $\pm$ 1.3 5.0 $\pm$ 1.2
@@
Average Feedback Length (words) 9.0 $\pm$ 6.1 9.1 $\pm$ 5.0 6.3 $\pm$ 3.8 6.6 $\pm$ 4.0
@@ 3 Fitness Interactive Coaching Dataset and Benchmark
We now introduce QEVD including QEVD-FIT-300K and QEVD-FIT-COACH dataset and benchmark in detail. We begin with a detailed description of the QEVD-FIT-COACH benchmark followed by additional details of QEVD-FIT-300K and QEVD-FIT-COACH datasets.
@@ 3.1 QEVD-FIT-COACH Benchmark
The QEVD-FIT-COACH benchmark contains videos of participants performing a structured workout while receiving live feedback. These feedbacks may be corrective, affirmative, or informative depending on user activity, to improve their form and pacing as they follow the workout. A trimmed workout session with two exercises, jumping jacks followed by high knees, is shown in Fig. 1. The segment begins with affirmative feedback to acknowledge the participant starting the exercise. Next, a series of feedbacks to correct user mistakes and affirm their compliance are provided. At first the participant exhibits low range of motion and is asked to correct this. Once the range of motion has improved, the participant receives an encouraging feedback. However, they then stop moving their arms and hence another corrective feedback is delivered. Note, the feedback accounts for them previously having moved their arms. The user then moves on to the high knees exercise after being requested to do so. Here, the user receives corrective feedback to raise their knees to the appropriate height and to improve their pace. These examples highlight the highly interactive coaching sessions in our QEVD-FIT-COACH benchmark and showcase the tight coupling between the participant actions and the timely feedback.
@@
Figure 2: Example annotations available on the short video clips from the QEVD-FIT-300K dataset. Annotations include question/answer pairs from our fitness questions dataset and feedbacks from a coaching perspective.
@@ 3.2 Fitness Datasets for Training
Together, the QEVD-FIT-300K and QEVD-FIT-COACH datasets in QEVD are designed to both instill domain understanding for fitness coaching and for providing effective feedbacks during live coaching sessions. It consists of three annotation types (Fig. 2) which are described in detail below.
@@
Fine-grained Fitness Activity Labels. This includes $460+$ hours of labeled (short) videos (QEVD-FIT-300K) crowd-sourced from over $1900$ unique participants in the wild. They cover $148$ different exercises and their variations including: varied pacing, performing common mistakes, and modified form. Exercise variations were determined top-down through consultation with expert fitness instructors. Participants were then provided detailed instructions and an accompanying reference video to perform the exercises and their pre-determined variations. Approx. $10$ variations were collected per exercise. We show such variations for the push-ups exercise in Fig. 2 (top). Additionally, there are $49$ types of general activities such as “grabbing a towel” or “drinking from a bottle”. Video lengths are in the $2-10$ second range. There are approximately $3,500$ videos per exercise on average ($300$k+ clips overall) and a total of $1,800$+ fine-grained classes capturing the exercise variations and general activities. These labels support training vision models for fine-grained understanding of human motion associated with fitness exercises.
@@
Fitness questions. In addition to the fine-grained labels, question-answer pairs querying video properties are provided for each fine-grained short video. The questions can be broadly divided into two types: high-level and fine-grained, as shown in Fig. 2. The high-level questions are directed at the overall exercise type and performance of the participant, e.g., in Fig. 2, high-level questions include “What exercise is the user doing?”, “Is the user doing this right?”. Fine-grained questions are designed to teach fine-grained details of exercises performed by the participant, e.g., “Was the user’s squat shallow and used only one arm?”, or “Is the user going as fast as possible?”. Overall statistics for this subset, including a breakdown of the provided splits, can be found in Tab. 2.
@@
Fitness feedbacks. We provide a set of annotations from a second-person perspective to support feedback in a live coaching session. It contains both fine-grained short videos mentioned above (QEVD-FIT-300K) and additional long-range videos (QEVD-FIT-COACH). For the fine-grained short videos, an average of $2$ feedbacks per video are provided (Tab. 2). These feedbacks occur at the end of each video, as shown in Tab. 2 (bottom), where the feedbacks focus on improving the form of the participant, specifically encouraging them to go deeper on the squats and to punch with both arms. We collect an additional ${\sim}9$ hours of fitness coaching sessions following the same methodology as the QEVD-FIT-COACH benchmark. They contain average of $5$ feedbacks per exercise (approximately $35$ per workout including instructions, Tab. 2).
@@
Figure 3: Architecture of the Stream-VLM model. The visual stream is processed by a 3D CNN, the language backbone is a LLaMA-2-7B model and the special action tokens (<next> and <feedback>) are highlighted in orange.

### v2

@@ Abstract
Vision-language models have shown impressive progress in recent years. However, existing models are largely limited to turn-based interactions, where each turn must be stepped (i.e., prompted) by the user. Open-ended, asynchronous interactions, where an AI model may proactively deliver timely responses or feedback based on the unfolding situation in real-time, are an open challenge. In this work, we present the Qevd benchmark and dataset, which explores human-AI interaction in the challenging, yet controlled, real-world domain of fitness coaching – a task which intrinsically requires monitoring live user activity and providing immediate feedback. The benchmark requires vision-language models to recognize complex human actions, identify possible mistakes, and provide appropriate feedback in real-time. Our experiments reveal the limitations of existing state-of-the-art vision-language models for such asynchronous situated interactions. Motivated by this, we propose a simple end-to-end streaming baseline that can respond asynchronously to human actions with appropriate feedback at the appropriate time.
@@ 1 Introduction
Datasets that combine visual information and language have greatly contributed to advancing the abilities of AI models over the past years, ranging from captioning [19], to visual questions answering [6], to visual dialogue [17], and beyond. Particularly impressive showcases of this progress are recent models such as GPT-4o [47] and Gemini [57], which can interact with users in real-time.
@@
Despite the impressive recent progress, existing vision-language models still lag far behind human capabilities. While state-of-the-art models can be queried (e.g., through prompting) to comment on events shown in the camera stream, they lack the ability to interact asynchronously with the user as demanded by the situation, rather than only when prompted. Such interactive scenarios that are grounded in the spatial and temporal context of an unfolding situation are commonly referred to as “situated” [5, 12, 13]. Addressing such situated interactive scenarios will be a key to developing real-world assistive vision-language models.
@@
A notable type of situated interaction is the instructional or coaching scenario, where an instructor guides a user through a complex activity, such as live fitness coaching. The real-world domain of live fitness coaching has several benefits that make it an ideal test-bed for studying situated assistive vision-language models: Firstly, fitness routines have a controlled structure in that users are expected to, but may not, follow a prescribed series of actions. Secondly, despite the structured nature of fitness routines, coaching in this domain remains a complex and unsolved problem for current vision-language systems. The nuances of human motion pose a significant challenge for these systems to effectively understand the dynamic situation and respond interactively. Finally, live fitness coaching represents a rapidly growing real-world application. The increasing popularity of home workouts [2], highlights the practical need and potential impact of developing effective solutions in this area. A successful vision-language model for live fitness coaching could thus offer users tangible benefits.
@@
Currently available large-scale video datasets [23, 26, 45, 54, 71] provide a rich set of annotations with expert demonstrations in domains such as cooking or house-hold activities. Expert demonstrations alone are not sufficient for real-world instructional scenarios, such as live fitness coaching, where the user may make mistakes. This has been addressed recently for some ego-centric instructional tasks by [10] and [61]. Successfully guiding a user through a fitness routine additionally necessitates the ability to understand fine-grained human actions, to provide appropriate instructions, and to provide corrective feedback—grounded in those fine-grained human actions—to correct any mistakes made by the user.
@@
Overall, our main contributions are: (1) We propose the first large scale benchmark and dataset, Qualcomm Exercise Videos Dataset (Qevd), aimed at the development of video-language models for live coaching. Qevd***Data and code are available at https://www.qualcomm.com/developer/software/qevd-dataset and https://github.com/Qualcomm-AI-research/FitCoach contains over 474+ hours of videos for fitness activity recognition and coaching. It includes short-clip videos (Qevd-Fit-300K) ($\sim$5 seconds in length) annotated with $1$M+ question-answer pairs, and long-range videos (>3 minutes in length) annotated with live feedback (Qevd-Fit-Coach, cf., Fig. 1); (2) We perform a comprehensive evaluation of state-of-the-art VLMs on the Qevd-Fit-Coach benchmark, revealing that the task is largely unsolved and offers significant room for improvement; (3) As a step towards closing the gap towards situated interaction, we propose a novel video-language model, Stream-VLM, which, instead of being limited to turn-based interactions, can decide on-the-fly when and what to say to the user. The model is trained end-to-end to perform real-time visual interaction with a user based on live camera input.
@@ 2 Related Work
Datasets for Activity Recognition. There is a large body of work on visual activity recognition. This includes NTU RGB+D 120 [50], FineGym [51], UCF101 [55], Kinetics [14], Moments in Time [46], ActivityNet [26], AVA-Kinetics [32], Charades [54], Something-Something [21], Something-Else [43], and others. Unlike our Qevd benchmark and dataset, these datasets do not contain detailed multi-modal annotations such as questions and feedbacks, or long videos with multiple actions or events. This restricts their utility in the development of interactive video-language models. Furthermore, while these datasets focus on a wide range of human activities, they contain only a few fitness activity related classes, if any (Tab. 1, col 2).
@@
Datasets for Procedural Activities. The Epic-Kitchens [16] dataset provides ego-centric videos of non-scripted daily kitchen activities, with post-hoc recorded narrations. YouCook2 [71] provides cooking videos annotated with instructions on how to prepare specific meals, largely featuring expert chefs. HowTo100M [45] provides narrated instructional videos, featuring a variety of activities. These datasets are not interactive (Tab. 1, col 4) as they feature first-person instructions or narrations. Furthermore, they largely feature experts and do not include mistakes likely to be made by novices (Tab. 1, col 5). In contrast, Qevd is interactive, features participants with diverse skill levels, and thereby, includes mistakes likely to be made by novices.
@@
Ego-Exo4D [23] includes a highly diverse set of activities, performed by participants with a variety of skill levels. Assembly-101 [49] features videos of people with diverse skill levels assembling and disassembling 101 “take-apart” toy vehicles. However, these datasets are not interactive and they do not include corrective feedbacks (Tab. 1, col 6). Qevd is interactive and includes corrective feedbacks from the perceptive of the fitness coach.
@@
Similar to our work, WTAG [10] and HoloAssist [61] include corrective feedbacks and are focused on the development of interactive AI assistants. However, they focus on domains such as cooking or object manipulation. As they are recorded from an ego-centric perspective, they do not contain complex human actions (Tab. 1, col 3). Furthermore, while they include mistakes and associated corrective feedbacks, they do not include diverse examples of possible mistakes per target task (Tab. 1, col 7). Qevd is recorded from the perspective of a virtual fitness coach and includes fine-grained human actions – diverse exercises and their variations including a wide diversity of mistakes per exercise.
@@
Datasets for VQA and Reasoning. ActivityNet Captions [26], Vatex [60], TRECVID [7], HD-VILA [63], TGIF [34], WebVid [8], Charades [54], STAR [62] and Agqa [24] among others, focus largely on video captioning and question answering tasks. Finally, there exist a wide range of datasets on visual reasoning [6, 29, 48], including visual dialogue, for example [17], all of which, in contrast to our work are based on still images rather than videos. FixMyPose [30] contains annotated instructions for pose correction but is limited to pairwise images and is activity-agnostic.
@@
Models for Situated Interactions. There is also a growing body of work on enabling LMs to generally reason over visual input [4, 25, 41, 53, 56, 59, 64, 66, 69]. However, the existing models can answer only high-level questions about depicted scenes and objects. Models for exercise feedback are discussed in [20, 52] among others. However, such models are based on the recognition of whether or not an exercise was performed, or on counting repetitions, rather than providing interactive guidance and reasoning about the user’s movements from a third-person viewpoint, which is the focus of this work.
@@
Total High-level Questions 535,299 31,326 - -
Total Fine-grained Questions 377,678 28,849 - -
@@
Total Feedbacks 573,637 36,333 5,403 2,484
Average Feedbacks per Video†† 2.0 $\pm$ 10.1 2.1 $\pm$ 10.2 5.0 $\pm$ 1.3 5.0 $\pm$ 1.2
@@
Average Feedback Length (words) 8.9 $\pm$ 5.1 9.2 $\pm$ 5.1 6.3 $\pm$ 3.8 6.6 $\pm$ 4.0
@@ 3 Fitness Interactive Coaching Dataset and Benchmark
We now introduce Qevd, including Qevd-Fit-300K, and the Qevd-Fit-Coach dataset and benchmark, in detail. We begin with a detailed description of the Qevd-Fit-Coach benchmark, followed by additional details of the Qevd-Fit-300K and Qevd-Fit-Coach datasets.
@@ 3.1 Qevd-Fit-Coach Benchmark
The Qevd-Fit-Coach benchmark contains videos of participants performing a structured workout while receiving live feedback. These feedbacks may be corrective, affirmative, or informative, depending on user activity, to improve their form and pacing as they follow the workout using the temporal structure described below.
Feedback Structure. The feedbacks in the Qevd-Fit-Coach benchmark have the following structure: At the start of each exercise, acknowledging feedback is given once the user has started; otherwise, a reminder to do so is provided. A corrective feedback is provided as soon as a mistake is clearly visible. Similarly, when the user begins to correct their mistake, feedback is provided to acknowledge and guide the user to successfully correct the error. If the user is performing the exercise correctly, feedback focuses on repetition counting. When repetition counts are not possible, such as with deltoid stretches, users receive positive, encouraging feedback, regularly with an average silence period of $5$ seconds between successive feedback. Finally, at the end of each exercise, a feedback focused on the overall performance during that exercise is provided. This temporal structure ensures that feedbacks in the Qevd-Fit-Coach benchmark occur at predictable time-steps, aligned to visually salient moments. The annotated feedbacks of each video were verified by a second annotator.
An example from the Qevd-Fit-Coach benchmark, illustrating a trimmed workout session with two exercises—jumping jacks followed by high knees—is shown in Fig. 1. The segment begins with affirmative feedback to acknowledge the participant starting the exercise. Next, a series of feedbacks to correct user mistakes and affirm their compliance are provided. Initially, the participant exhibits a low range of motion and is asked to correct this. Once the range of motion improves, the participant receives encouraging feedback. However, they then stop moving their arms, incurring another corrective feedback. Note that this feedback considers their previous arm movements. The user then moves on to the high knees exercise after being requested to do so. Here, the user receives corrective feedback to raise their knees to the appropriate height and to improve their pace. The session ends with positive encouraging feedback acknowledging that the user has currently performed the exercise. These examples highlight the highly interactive coaching sessions in our Qevd-Fit-Coach benchmark and showcase the tight coupling between the participant actions and timely feedback.
@@
Figure 2: Example annotations available on the short video clips from the Qevd-Fit-300K dataset. Annotations include question/answer pairs from our fitness questions dataset and feedback from a coaching perspective.
@@ 3.2 Fitness Datasets for Training
Together, the Qevd-Fit-300K and Qevd-Fit-Coach datasets in Qevd are designed to instill domain understanding for fitness coaching and provide effective feedbacks during live coaching sessions. They consist of three annotation types (Fig. 2), which are described in detail below.
@@
Fine-grained Fitness Activity Labels. This includes $460+$ hours of labeled (short) videos (Qevd-Fit-300K) crowd-sourced from over $1,900$ unique participants in the wild. They cover $148$ different exercises and their variations including: varied pacing, performing common mistakes, and modified form. Exercise variations were determined top-down through consultation with expert fitness instructors. Participants were then provided detailed instructions and an accompanying reference video to perform the exercises and their pre-determined variations. Approximately $10$ variations were collected per exercise. We show such variations for the push-ups exercise in Fig. 2 (top). Additionally, there are $49$ types of general activities such as “grabbing a towel” or “drinking from a bottle”. Video lengths are in the $2$ to $10$ second range. There are approximately $3,500$ videos per exercise on average (${\sim}300$k clips overall) and a total of $1,800$+ fine-grained classes capturing the exercise variations and general activities. Each video and corresponding labels were manually reviewed for correctness by at least one unique crowd-worker. These labels support training vision models for fine-grained understanding of human motion associated with fitness exercises.
@@
Fitness Questions. In addition to the fine-grained labels, question-answer pairs querying video properties are provided for each fine-grained short video. This data can be used to provide further grounding of the LM’s concepts in the observed visual input. The questions can be broadly divided into two types: high-level and fine-grained, as shown in Fig. 2. The high-level questions are directed at the overall exercise type and performance of the participant, e.g., in Fig. 2, high-level questions include “What exercise is the user doing?”, “Is the user doing this right?”. Fine-grained questions are designed to teach fine-grained details of exercises performed by the participant, e.g., “Was the user’s squat shallow and used only one arm?”, or “Is the user going as fast as possible?”. These questions are generated using the Mixtral-8B-Instruct LLM [28] and the question generation process follows the scheme detailed in [37]. Overall statistics for this subset, including a breakdown of the provided splits, can be found in Tab. 2.
@@
Fitness Feedbacks. We provide a set of annotations from a second-person perspective to support feedback in a live coaching session. It contains both fine-grained short videos mentioned above (Qevd-Fit-300K) and additional long-range videos (Qevd-Fit-Coach). For the fine-grained short videos, an average of $2$ feedbacks per video are provided (Tab. 2). These feedbacks occur at the end of each video, as shown in Fig. 2 (bottom). In the shown example, the feedbacks focus on improving the form of the participant, specifically, by encouraging them to squat deeper and punch with both arms. We collected an additional ${\sim}9$ hours of fitness coaching sessions following the same methodology as the Qevd-Fit-Coach benchmark. These sessions contain an average of $5$ feedbacks per exercise, totaling approximately $35$ feedbacks per workout, including instructions (See Tab. 2).
@@
Figure 3: Architecture of the Stream-VLM model. The visual stream is processed by a 3D CNN and the language backbone is a LLaMA-2-7B model; special action tokens (<next> and <feedback> ) are highlighted in orange.